OpenStudy (anonymous):

what is the domain of the composite function G(F(x))? given F(x)=3-x G(y)=√y g(F(x))=√3-x

OpenStudy (mathstudent55):

g(f(x)) = sqrt(3 - x) What values are not allowed inside a square root?

OpenStudy (anonymous):

this are the answers to it x<0 x>3 x≤3 x<3 x≤0 all real numbers

OpenStudy (mathstudent55):

Do you want to understand how to solve this and other similar problems, or you just want the solution?

OpenStudy (anonymous):

i want to know how to solve it

OpenStudy (mathstudent55):

Then look above and answer my question.

OpenStudy (anonymous):

negatives right?

OpenStudy (mathstudent55):

Good. Inside a square root sign, you can't have negatives. Tha means you can have zero and positive.

OpenStudy (mathstudent55):

Inside the square root symbol, you have 3 - x. 3 - x can be zero or positive. that means you have the inequality 3 - x >= 0

OpenStudy (mathstudent55):

Now you solve that inequality for x to get the answer.

OpenStudy (anonymous):

so then you subtract 3 right?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

so the answer would be x>3?

OpenStudy (mathstudent55):

3 - x >= 0 -x >= -3 Divide both sides by -1, but remember to change direction of inequality sign: x <= 3

OpenStudy (anonymous):

thats when you divide only by the negative right?

OpenStudy (mathstudent55):

Correct, but it's true if you multiply or divide by a negative number.

OpenStudy (anonymous):

ooohh okay i see